the width of a rectangle is 9cm less than it's length.If the area is 190 square cm. Find the length and width of the rectangle.
Found 2 solutions by ikleyn, 68271:
Answer by ikleyn(52781)

.
19 cm and 10 cm.


Answer by 68271(1)

Given Area of Rectangle as Length(L) x W (Width)
L = x
W = (x - 9cm)
Area = 190 square cm
190 = x(x - 9)
190 = x^2 - 9x
x^2 - 9x = 190
x^2 - 9x - 190 = 0
x^2 - 19x - 10x - 190 = 0
x(x - 19) - 10(x - 19) = 0
(x - 19) (x - 10) = 0
x = 19, x = 10
Length = 19
Width = 10